Trace the following curves:


i) 𝑥(4𝑎


2 + 𝑦


2


) = 8𝑎


3


ii) 𝑦


2


(𝑥


2 + 𝑦


2


) + 𝑎


2


(𝑥


2 − 𝑦


2


) = 0


iii) 𝑎


2𝑦


2 = 𝑥


2


(𝑎


2 − 𝑥


2


)


iv) 𝑥 = 𝑡


2


, 𝑦 = 𝑡 −


𝑡


3


3



v) 𝑟 = 𝑎(1 + 𝑠𝑖𝑛𝜃)


vi) 𝑟


2 = 𝑎


2


𝑠𝑖𝑛2𝜃


vii) 𝑟 = 𝑎 𝑐𝑜𝑠3𝜃


viii) 𝑟 = 𝑎(1 + 𝑠𝑖𝑛𝜃)

1
Expert's answer
2022-03-07T17:04:05-0500

A )4a²sq. units

B) 158a² sq. units

C

9
16a2​ sq. units

D

none of these


Correct option is B)

ay²=x²(a−x)⇒y=± xa
a−x​Curve tracing : y=xaa−x​We must have x≤aFor 0<x≤a, y>0 and for x<0,y<0Also y=0⇒x=0,aCurve is symmetrical about x-axis.When x→−∞,y→−∞Also, it can be verified that y has only one point of maxima for 0<x<a.Area =2∫0a​xaa−x​dxaa−x​=t⇒−ax=t²⇒x=a(1−t2)⇒A=2∫10a(1−t²)t(−2at)dt=41​(t²−t⁴)dt=4a²[3t³−5t⁵]01=4a²1]=158a²sq. units